# HateCOT: An Explanation-Enhanced Dataset for Generalizable Offensive Speech Detection via Large Language Models

### v1

@@ 5.2 QA 2: Inter-model Comparison
Using the template in Figure 6, we collect numerical ratings from GPT-4 on a scale from 1 (least) to 5 (very) on these criteria for 200 posts per model per dataset, for a total of 2,400 samples. Figure 3 displays the mean ratings for each model-dataset pair, as well as Overall scores averaged across all datasets for each model. Ratings for Informativeness appear consistently lower than those of the other 2 criteria, likely due to the challenging context-dependent nature of generating post-specific relevant information. Overall, Persuasiveness and Soudness ratings around 4 out of 5 suggest high likelihood of acceptance by human users for content moderation.
We also note the general correlation between the models’ classification performance (Figure 1(b)) and mean ratings on each dataset. All finetuned models attain F-1 scores above 0.9 along with the highest average ratings on HateCheck, followed by HateXplain. Latent_Hate and its more difficult subset Implicit_Hate yield markedly lower average ratings. We perform ANOVA test on the ratings of models stratified by dataset and overall using $\alpha$=0.05. Aggregated ratings do not appear to differ meaningfully across criteria. Interestingly, COT-T5-XL’s superior performance over Llama models on Implicit_Hate coincides with statistically or marginally significant differences between average ratings. This correlation between better classification and higher quality of explanation is useful to calibrate models for real-life deployment.
5.3 QA 3: Target Identification
Dataset COT-T5-XL Llama 7B Llama 13B
HateCheck 94.5 % 96.5 % 93.5 %
HateXplain 71.0 % 74.0 % 74.0 %
Latent_Hate 50.5 % 51.0 % 44.5 %
Finally, we investigate the in-domain finetuned models’ capabilities to identify the target of the sentiments expressed by the post. We also randomly select 200 samples from each dataset that have Target variable annotated by humans, then ask GPT-4 to judge whether the explanations from the models mention at least one of the listed targeted groups using the template in Figure 7. Note that we encourage GPT-4 to consider potential variance of expression and not restrict to exact matches.
@@ 7 Limitations
We first acknowledge that our work is restricted to English corpora. Second, due to computational limitations, we could not perform experiments on larger open-source models. With the development of newer, more powerful models, it is reasonable to expect their performance to further improve though the use of this dataset of our corpus. Finally, an elaborated human study may further corroborate our quality analysis of explanations. Given supporting literature , we expect our findings to remain consistent Gilardi et al. (2023); Yang et al. (2023); Lin et al. (2023).
@@ 8 Ethics Statement
The authors note the risk of propagating implicit biases that LLMs are known to carry Cheng et al. (2023); Gupta et al. (2023). We do acknowledge the potential malicious usage of our corpus to generate content capable of evading detection, or jeopardizing classifiers’ performance.
@@ A.2.1 Inferencing
We used OpenAI’s API to access the publicly available version of GPT-3.5 in November 2023, and GPT-4 in January 2024. To obtain explanations from the former (as described in 3.2.2), we set the temperature to 0.7 during inference to obtain multiple outputs per sample. For GPT-4 and other open-source models, we use greedy decoding.
@@ A.3 Variants of Prompt Template for Explanation
Model Base Pretrained
COT-T5-XL
11 (22%) 39 (78%)
Llama 7B
20 (40%) 30 (60%)
Llama 13B
23 (46%) 27 (54%)
Table 5: Number of times human annotator prefers outputs of Base vs Pretrained in Section 5.1
@@ A.4 Human Annotation for QA Experiment
With approved IRB, we recruit a native English speaker from the United States using the Prolific platform to annotate 50 samples per model, for a total of 150 data points for the task described in Section 5.1. The annotator was paid a fair wage at \$15 per hour, and forewarned about the nature of the task. Table 5 shows the results. We observe that preference for explanations generated by Pretrained models remains consistent with GPT-4’s.
@@
Neutral : A post that is not offensive to any group of people. Hateful: An offensive post, motivated, in whole or part, by the writer’s bias against an aspect of a group of people.
Qian et al.,2019
@@
Not Hate : Does not contain any attack on people based on protected characteristics. Hate: Contains direct attack on people based on protected characteristics such as race, ethnicity, national origin, religious affiliation, sexual orientation, caste, sex, gender, gender identity, and serious disease or disability.
Sap et al.,2019
@@
Not Offensive : not offensive to anyone. Offensive : denotes the overall rudeness, disrespect, or toxicity of a post. whether a post could be considered offensive to anyone.
Vidgen et al.,2021a
@@
Neutral: Content that does not fall into other categories, usually entirely unrelated to abuse, hate, prejudice, or intolerance. Identity Directed Abuse: Content that directs abuse at an identity, which relates to fundamental aspects of individuals’ social position, community and self-representation. An identity includes but is not limited to religion, race, ethnicity, gender, sexuality and sexual preferences, immigration status, nationality, ableness, physical appearance and class. Affiliation Directed Abuse : Content that directs abuse at people who have a voluntary affiliation with a profession, membership, association, ideology, or other well-defined group or collective. Person Directed Abuse : Content that directs abuse at an identifiable person.
Vidgen et al.,2021b
@@
Derogation : content which explicitly attacks, demonizes, demeans or insults a group. Animosity : content which expresses abuse against a group in an implicit or subtle manner. Threatening : content which expresses intention to, support for, or encourages inflicting harm on a group, or identified members of the group. Support For Hateful Entities : content which explicitly glorifies, justifies or supports hateful actions, events, organizations, tropes and individuals collectively, entities. Dehumanization : content which perceives or treats people as less than human. Not Hate : content that falls into none of the other categories.
Basile et al.,2019
@@
Hate Speech : language that is used to expresses hatred towards a targeted group or is intended to be derogatory, to humiliate,or to insult the members of the group. may also be language that threatens or incites violence. Offensive Language : may contain offensive terms but targets disadvantaged social groups in a manner that is potentially harmful to them. Neither : language that does not all into either of the other categories.
Hartvigsen et al.,2022
@@
Benign: Text that is not harmful nor offensive to anyone. Toxic: Text that could be seen as harmful to anyone, may contain offensive, rude, humor, insults, personal attacks, profanity, aggression, may refer to targeted group with harmful intent that is expressed in stereotypes or lewd manners
Toraman et al.,2022
@@
Hate : target, incite violence against, threaten, or call for physical damage for an individual or a group of people because of some identifying trait or characteristic. Offensive : humiliate, taunt, discriminate, or insult an individual or a group of people in any form, including textual. Normal : does not fall into any of the other categories.
@@
Table 6: Labels and definitions of 8 datasets used to synthesize training corpus.
@@
HateCheck
Non-hateful: Language that does not contain any abusive sentiment targeted at a protected group. Hateful: Abuse that is targeted at a protected group or at its members being a part of that group. Protected groups include age, disability, gender identity, familial status, pregnancy, race, national or ethnic origins, religion, sex or sexual orientation.
HateXplain
Normal : neither hate speech nor offensive. Hate Speech : language which attacks, demeans, offends, threatens, or insults a group based on race, ethnic origin, religion, disability, gender, age, sexual orientation, or other traits. it is not the presence of certain words that makes the text hate speech, rather you should look the context the word is used in the text. Offensive Language : usage of rude, hurtful, derogatory, obscene or insulting language to upset or embarrass people.
Latent_Hate
Not Hate : speech or actions that do not involve any form of hatred, prejudice, or discrimination toward individuals or groups based on their characteristics. Explicit Hate: openly expressed, direct forms of hatred and prejudice toward individuals or groups based on their characteristics. Implicit Hate: coded or indirect language that disparages a person or group on the basis of protected characteristics like race, gender, and cultural identity.
Implicit_hate
White Grievance : includes frustration over a minority group s perceived privilege and casting majority groups as the real victims of racism. this language is linked to extremist behavior and support for violence. Incitement To Violence : includes flaunting in group unity and power or elevating known hate groups and ideologies. Inferiority Language : implies one group or individual is inferior to another, and it can include dehumanization denial of a person s humanity , and toxic language that compares the target with disease, insects, animals . related to assaults on human dignity, dominance, and declarations of superiority of the in group. Irony : refers to the use of sarcasm , humor, and satire to attack or demean a protected class or individual. Stereotypes And Misinformation : associate a protected class with negative attributes such as crime, or terrorism. includes misinformation that feeds stereotypes and vice versa, like holocaust denial and other forms of historical negationism. Threatening And Intimidation : conveys a speaker’s commitment to a target s pain, injury, damage, loss, or violation of rights, threats related to implicit violation of rights and freedoms, removal of opportunities, and more subtle forms of intimidation.

### v2

@@ 5.2 QA 2: Inter-model Comparison
Using the template in Figure 7, we collect numerical ratings from GPT-4 on a scale from 1 (least) to 5 (very) on these criteria for 200 posts per model per dataset, for a total of 2,400 samples. Figure 3 displays the mean ratings for each model-dataset pair, as well as Overall scores averaged across all datasets for each model. Ratings for Informativeness appear consistently lower than those of the other criteria, likely due to the challenging context-dependent nature of generating post-specific relevant information. Overall, Persuasiveness and Soudness ratings around 4 out of 5 suggest high likelihood of acceptance by human users for content moderation. We also note the general correlation between the models’ better classification performance (Figure 2(b)) and higher mean ratings on each dataset, a useful artifact to calibrate models.
@@ 7 Limitations
We first acknowledge that our work is restricted to English corpora, a common limitation among literature on offensive speech Yin and Zubiaga (2021); Poletto et al. (2021). However, our approach sets a proof-of-concept for researchers to construct similar corpus in other languages by leveraging existing resources. Furthermore, our developmental pipeline is considerably more data-efficient than conventional approaches (Section 4.1, 4.2), potentially lowering the barrier of entry for practitioners without access to abundant resources. Therefore, this work invites further expansion on multilingual datasets, particularly to develop corpora with clearly defined definitions to facilitate synergy with other research.
Second, due to computational limitations, we could not perform experiments on larger open-source models. With the development of newer, more powerful models, it is reasonable to expect their performance to further improve though the use of this dataset of our corpus, as demonstrated by our empirical results.
Finally, we recognize the risk of propagating implicit biases that LLMs are known to carry Cheng et al. (2023); Gupta et al. (2023). However, we note that the approach of using LLMs (GPT-3.5-Turbo in this paper) to bridge the logical gaps in original rationales has been shown to produce outputs less prone to logical failures Yang et al. (2023). Biases in Pretrained Language Models have been attracting much attention in the research community. We invite further works to consider our approach to reduce hallucinations and biases in text generation.
@@ 8 Ethics Statement
We acknowledge the potential malicious usage of our corpus to generate content capable of evading detection, or jeopardizing classifiers’ performance.
@@ A.1 Data Pre-Processing
Dataset Best Model @ K =256 F1 Retrained F1 Base + Full Retrained F1 % Retrained Data Size Full Training Data Size Retrained Data %
HateCheck LLAMA 13B 0.95 0.99 96% 512 1,864 27%
HateXplain LLAMA 13B 0.64 0.72 89% 768 12,088 6%
Latent_Hate LLAMA 13B 0.66 0.64 103% 768 11,460 7%
Implicit_Hate COT-T5-XL 0.56 0.38 147% 1,536 2,707 57%
Table 4: Comparison of performance metrics for the best performing model at K=256 in-domain finetuning vs. finetuned on the full training set, without explanations or definitions. Retrained F1% denotes the percentage of macro F-1 score on the K=256 retrained model over that of the version trained on full data. Similarly, Retrained Data % denotes the percentage of data size used by the Retrained model over the size of the full data.
@@ A.2.1 Inferencing
We used OpenAI’s API to access the publicly available version of GPT-3.5 in November 2023, and GPT-4 in January 2024. To obtain explanations from the former (as described in 3.2.2), we choose temperature among candidates $\{0.3,0.5,\textbf{0.7}\}$ and settle on the last value during inference. This value is selected based on literature and multiple iterations of qualitative analysis of outputs Yang et al. (2023); Kim et al. (2023). For GPT-4 and other open-source models, we use greedy decoding.
@@ A.3 Variants of Prompt Template for Explanation
Dataset COT-T5-XL Llama 7B Llama 13B
HateCheck 94.5 % 96.5 % 93.5 %
HateXplain 71.0 % 74.0 % 74.0 %
Latent_Hate 50.5 % 51.0 % 44.5 %
@@ A.4 Human Annotation for QA Experiment
With approved IRB, we recruit 13 crowdsource workers using the Amazon Mechanical Turk platform to annotate 50 samples per model, for a total of 150 data points for the task described in Section 5.1. The annotators was paid a fair wage at \$15 per hour, and forewarned about the nature of the task. Annotators must be fluent English speakers. We also limit each annotator to no more than 100 posts (60% of the total 150 samples per model) to maintain diversity of opinions. We observe that preference for explanations generated by Pretrained models remains consistent with GPT-4’s.
The general demographic breakdown of the 13 annotators are described below:
  • Gender: Female (8), Male (5)
  • Age: 18-29 (2), 30-39 (4), 40-49 (4), 50+ (3)
  • Education: High School (2), 2-year college (5), 4-year college (4), Master’s or Higher (2)
A.5 QA 3: Target Identification
Finally, we investigate the in-domain finetuned models’ capabilities to identify the target of the sentiments expressed by the post. We also randomly select 200 samples from each dataset that have the Target variable annotated by humans, then ask GPT-4 to judge whether the explanations from the models mention at least one of the listed targeted groups using the template in Figure 8. Note that we encourage GPT-4 to consider potential variance of expression and not restrict to exact matches.
@@ A.5 QA 3: Target Identification
Salminen et al.,2018 Neutral : A post that is not offensive to any group of people. Hateful: An offensive post, motivated, in whole or part, by the writer’s bias against an aspect of a group of people.
@@
Qian et al.,2019 Not Hate : Does not contain any attack on people based on protected characteristics. Hate: Contains direct attack on people based on protected characteristics such as race, ethnicity, national origin, religious affiliation, sexual orientation, caste, sex, gender, gender identity, and serious disease or disability.
@@
Sap et al.,2019 Not Offensive : not offensive to anyone. Offensive : denotes the overall rudeness, disrespect, or toxicity of a post. whether a post could be considered offensive to anyone.
@@
Vidgen et al.,2021a Neutral: Content that does not fall into other categories, usually entirely unrelated to abuse, hate, prejudice, or intolerance. Identity Directed Abuse: Content that directs abuse at an identity, which relates to fundamental aspects of individuals’ social position, community and self-representation. An identity includes but is not limited to religion, race, ethnicity, gender, sexuality and sexual preferences, immigration status, nationality, ableness, physical appearance and class. Affiliation Directed Abuse : Content that directs abuse at people who have a voluntary affiliation with a profession, membership, association, ideology, or other well-defined group or collective. Person Directed Abuse : Content that directs abuse at an identifiable person.
@@
Vidgen et al.,2021b Derogation : content which explicitly attacks, demonizes, demeans or insults a group. Animosity : content which expresses abuse against a group in an implicit or subtle manner. Threatening : content which expresses intention to, support for, or encourages inflicting harm on a group, or identified members of the group. Support For Hateful Entities : content which explicitly glorifies, justifies or supports hateful actions, events, organizations, tropes and individuals collectively, entities. Dehumanization : content which perceives or treats people as less than human. Not Hate : content that falls into none of the other categories.
@@
Basile et al.,2019 Hate Speech : language that is used to expresses hatred towards a targeted group or is intended to be derogatory, to humiliate,or to insult the members of the group. may also be language that threatens or incites violence. Offensive Language : may contain offensive terms but targets disadvantaged social groups in a manner that is potentially harmful to them. Neither : language that does not all into either of the other categories.
@@
Hartvigsen et al.,2022 Benign: Text that is not harmful nor offensive to anyone. Toxic: Text that could be seen as harmful to anyone, may contain offensive, rude humor, insults, personal attacks, profanity, aggression, may refer to targeted group with harmful intent that is expressed in stereotypes or lewd manners
@@
Toraman et al.,2022 Hate : target, incite violence against, threaten, or call for physical damage for an individual or a group of people because of some identifying trait or characteristic. Offensive : humiliate, taunt, discriminate, or insult an individual or a group of people in any form, including textual. Normal : does not fall into any of the other categories.
@@
Table 6: Labels and definitions of 8 datasets used to synthesize training corpus. Definitions are lifted directly from their original works, with exceptions of minor adjustment for stylistic consistency.
@@
HateCheck Non-hateful: Language that does not contain any abusive sentiment targeted at a protected group. Hateful: Abuse that is targeted at a protected group or at its members being a part of that group. Protected groups include age, disability, gender identity, familial status, pregnancy, race, national or ethnic origins, religion, sex or sexual orientation.
HateXplain Normal : neither hate speech nor offensive. Hate Speech : language which attacks, demeans, offends, threatens, or insults a group based on race, ethnic origin, religion, disability, gender, age, sexual orientation, or other traits. it is not the presence of certain words that makes the text hate speech, rather you should look the context the word is used in the text. Offensive Language : usage of rude, hurtful, derogatory, obscene or insulting language to upset or embarrass people.
Latent_Hate Not Hate : speech or actions that do not involve any form of hatred, prejudice, or discrimination toward individuals or groups based on their characteristics. Explicit Hate: openly expressed, direct forms of hatred and prejudice toward individuals or groups based on their characteristics. Implicit Hate: coded or indirect language that disparages a person or group on the basis of protected characteristics like race, gender, and cultural identity.
Implicit_hate White Grievance : includes frustration over a minority groups perceived privilege and casting majority groups as the real victims of racism. This language is linked to extremist behavior and support for violence. Incitement To Violence : includes flaunting in group unity and power or elevating known hate groups and ideologies. Inferiority Language : implies one group or individual is inferior to another, and it can include dehumanization, denial of a person’s humanity , and toxic language that compares the target with disease, insects, animals . Related to assaults on human dignity, dominance, and declarations of superiority of the in group. Irony : refers to the use of sarcasm , humor, and satire to attack or demean a protected class or individual. Stereotypes And Misinformation : associate a protected class with negative attributes such as crime, or terrorism. includes misinformation that feeds stereotypes and vice versa, like holocaust denial and other forms of historical negationism. Threatening And Intimidation : conveys a speaker’s commitment to a target’s pain, injury, damage, loss or violation of rights, threats related to implicit violation of rights and freedoms, removal of opportunities, and more subtle forms of intimidation.
@@
(a) Example from Toraman et al. (2022)
(b) Example from Salminen et al. (2018)
(c) Example from Sap et al. (2019)
Figure 4: Examples drawn from our training corpora showing their native Post, Target and Rationale, along with the corresponding GPT-3.5-Turbo-enhanced explanations. Due to their nature as fragmented annotations, verbatim Rationales are not serviceable explanation, but can serve as guiding signals that leverage GPT’s generative capabilities to construct legible passages with detailed justifications.